Calculate the total amount of energy released from the fission of 1 kg of U-235.
1
Expert's answer
2020-06-04T09:01:23-0400

 



"Q = (235.0439 - 141.9077- 89.9047 - 3\cdot1.0086) = 0.205503\ amu"



"Q = 191.4\ MeV = 3.15\cdot10^{-11} J"

"E=( 3.15\cdot10^{-11})\frac{ 6.023\cdot10^{23}}{0.235}= 8.06\cdot10^{13}J"